### v1

@@ 6.2 Numerical results for Cahn-Hilliard system
We proceed in reporting experimental studies to assess the accuracy of the proposed framework. Consider a cylinder revolving around the $z$-axis with radius $R=1$ and height $l=2$. To satisfy the assumption $\nabla_{\Gamma}\cdot{\boldsymbol{v}}=0$, we transform the initial cylinder through the isometry ${\boldsymbol{\Phi}}(t,{\boldsymbol{p}})={\boldsymbol{\Phi}}^{\mathcal{A}}(t,{\boldsymbol{p}})=\mathbb{A}(t){\boldsymbol{p}}+B(t)$
(58) $\mathbb{A}(t)=\begin{bmatrix}\lambda(t)&0&0\\ 0&\lambda(t)&0\\ 0&0&1/\lambda(t)\end{bmatrix},\quad B(t)=\begin{bmatrix}0.2t\\ 0.1t\\ 0\end{bmatrix}$,
@@
    chSolver 1: Solver 56 with no postprocessing.
@@
    chSolver 2: Solver 56 with imposed bound preservation in the interval (1,1)(-1,1).
@@
    chSolver 3: Solver 56 with imposed mass preservation using the lumped-mass inner product $(\cdot\,,\,\cdot)^{h}$.
@@
    chSolver 4: Solver 56 with imposed bound preservation in the interval (1,1)(-1,1) and mass preservation.
Zero flux Neumann boundary conditions are imposed for both $u$ and $w$. A random uniform initial distribution in the interval [1,1][-1,1] is chosen for $u_{0}$. The parameters are set as
@@ 7 Membrane force system
In the realm of mathematical modeling, principal curvatures have long been used for the theory of elastic plates and shells, with the work of Kirchhoff [66] representing what may be the most famous case. Later in the ’70s, the work of Canham [27] and Helfrich [54] on the link between biological membranes shape and curvature functionals has de facto initiated an entirely new field of research focused on the use of geometry of hypersurfaces for biophysical modeling.
@@
(59) $\mathcal{E}_{B}(t)=\int_{\Gamma}\frac{\gamma_{W}}{2}(\kappa-\kappa_{0})^{2}+\int_{\Gamma}\gamma_{G}\kappa_{G}+\gamma_{\partial}\int_{\partial\Gamma}1,$
where $\gamma_{W},\;\gamma_{G}\in\mathbb{R}$ are bending rigidities, $\kappa:\mathcal{G}_{T}\to\mathbb{R}$ is the previously introduced mean curvature, $\kappa_{0}\in\mathbb{R}$ is the spontaneous curvature, and $\kappa_{G}:\mathcal{G}_{T}\to\mathbb{R}$ is the Gaussian curvature $\kappa_{G}=\det\mathbb{W}$. The parameter $\gamma_{\partial}$ takes into account the possible line energy of $\partial\Gamma$. For the sake of simplicity we will consider the case $\gamma_{G}=constant$, $\gamma_{W}=constant$, and assume that no topological change occur along the evolution. Taken into account these simplifications, the gradient flow of a membrane $\Gamma$ is a family of evolving surfaces with normal velocity $\boldsymbol{v}$$^{\perp}$ given by [88, 14]
(60) ${\boldsymbol{v}}^{\perp}=\gamma_{W}\Bigl(-\Delta_{\Gamma}\kappa+\frac{1}{2}\kappa(\kappa-\kappa_{0})^{2}-(\kappa-\kappa_{0})|\mathbb{W}|^{2}\Bigr){\boldsymbol{n}}_{\Gamma}.$
@@
Remark 7.1.
Other geometrical flows like the mean curvature flow and the surface diffusion flow are also of interest in this field. Being somewhat simpler, extensive research is already available for them, while less is known about the Willmore flow, which we focus on in this article. We refer to [38, 32, 15] for further details.
Remark 7.2.
@@ 7.1 Numerical methods for Helfrich system
We choose here the family of BGN-type algorithms as pioneered in [11] and later adapted for Willmore flow in [12, 14]. The schemes pivot around the two following ingredients:
  • Introduce the mean curvature vector ${\boldsymbol{\kappa}}(t)=\kappa(t){\boldsymbol{n}}_{\Gamma}(t)$ as independent variable and use the defining equation for the mean curvature vector ${\boldsymbol{\kappa}}=\Delta_{\Gamma}\mathbf{Id}_{\Gamma}$ in order to reduce the fourth order problem to a coupled second order one [11];
  • Use the weak formulation to rewrite and simplify the term $\kappa|\mathbb{W}|^{2}$ (more precisely $\kappa|\mathbb{H}|^{2}$ in the immersed setting) taking advantage of the newly introduced variable $\boldsymbol{\kappa}$}. This reformulation not only simplifies and linearizes the system but also leads to stable schemes as proven in [37].
Our scheme choice builds on the one derived in [14] for boundary value problems. Among the different types of boundary conditions considered in [14] we will restrict to the case of clamped boundary conditions, for which
@@
(62a) $({\boldsymbol{v}}^{\perp}\,,\,{\boldsymbol{\phi}})=~(\nabla_{\Gamma}{\boldsymbol{y}}\,,\,\nabla_{\Gamma}{\boldsymbol{\phi}})+(\nabla_{\Gamma}\cdot{\boldsymbol{y}}\,,\,\nabla_{\Gamma}\cdot{\boldsymbol{\phi}})-((\nabla_{\Gamma}{\boldsymbol{y}})^{T}\,,\,\mathcal{D}({\boldsymbol{\phi}})\;\mathbb{P}_{\Gamma})$
@@
(62b) $({\boldsymbol{y}}\,,\,{\boldsymbol{\psi}})=~(\gamma_{W}({\boldsymbol{\kappa}}-\kappa_{0}{\boldsymbol{n}}_{\Gamma})\,,\,{\boldsymbol{\psi}}),$
(62c) $({\boldsymbol{\kappa}}\,,\,{\boldsymbol{\xi}})=~-(\mathbb{P}_{\Gamma}\,,\,\nabla_{\Gamma}{\boldsymbol{\xi}})+({\boldsymbol{\mu}}\,,\,{\boldsymbol{\xi}})_{\partial\Gamma},$
@@
(63) ${\boldsymbol{v}}_{h}^{\perp}(t_{n})=\frac{1}{\tau}\sum_{i}({\boldsymbol{y}}_{i}(t_{n+1})-{\boldsymbol{y}}_{i}(t_{n})){\boldsymbol{\phi}}_{i}(t_{n})=\frac{1}{\tau}{\boldsymbol{d}}_{h}^{n+1},$
where we recall that ${\boldsymbol{y}}_{i}$ are the mesh points and $\{{\boldsymbol{\phi}}_{i}\}$ the associated vector-valued basis such that ${\boldsymbol{\phi}}_{i}({\boldsymbol{y}}_{j})=\delta_{ij}$ at all times. Moreover, (62c) and (62b) can be collapsed to
@@
(64) $+(\kappa_{0}{\boldsymbol{n}}_{\Gamma^{n}}\,,\,{\boldsymbol{\psi}})_{\Gamma^{n}}+({\boldsymbol{\mu}}\,,\,{\boldsymbol{\psi}})_{\partial\Gamma^{n}}.$
@@
(65a) $\mathopen{}\left(\frac{{\boldsymbol{d}}_{h}^{n+1}}{\tau}\,,\,{\boldsymbol{\phi}}_{h}\mathclose{}\right)_{\Gamma_{h}^{n}}^{h}-(\nabla_{\Gamma_{h}^{n}}{\boldsymbol{y}}^{n+1}_{h}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\phi}}_{h})_{\Gamma_{h}^{n}}=~(\nabla_{\Gamma_{h}^{n}}\cdot{\boldsymbol{y}}^{n}_{h}\,,\,\nabla_{\Gamma_{h}^{n}}\cdot{\boldsymbol{\phi}}_{h})_{\Gamma_{h}^{n}}$
@@
(65b) $\frac{1}{\gamma_{W}}({\boldsymbol{y}}^{n+1}_{h}\,,\,{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}^{h}+(\nabla_{\Gamma_{h}^{n}}{\boldsymbol{d}}_{h}^{n+1}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}$
@@
Remark 7.4.
Note that (65) corresponds to the scheme presented in [12] for the choice of the parameter $\theta=1$, i.e. without tangential motion control.
@@ 7.2 Numerical results for Helfrich system
To test the accuracy of our method for gradient flows, we start by performing classical studies on Willmore flow such as the ones in [37, 12, 14]. For every test we compare
@@
    hfSolver 2: algorithm (65) together with the mesh redistribution described in (33). We highlight that it is of crucial importance that, when the algorithm in Section 4 is employed, geometrical quantities characteristic of the mesh are updated. In our case, we refer to the mean curvature ${\boldsymbol{\kappa}}_{h}$ and the auxiliary variable ${\boldsymbol{y}}_{h}$ of the algorithm in (65).
We start by testing the convergence using the example presented in [12, Section 5.3]. It is an evolving sphere with initial radius $R_{0}$ and spontaneous curvature $\kappa_{0}$. The radius evolution $R(t)$ is the solution to the ordinary differential equation
(66) $R_{t}=-\frac{\kappa_{0}}{R}\Bigl(\frac{2}{R}+\kappa_{0}\Bigr),\quad R(0)=R_{0}>0.$
@@
(67) $\lVert\Gamma-\Gamma_{h}\rVert_{L^{\infty}(L^{\infty})}=\max_{n}\max_{i}|\lVert{\boldsymbol{y}}_{i}(t_{n})\rVert-R(t_{n})|,$
@@
Figure 8: Convergence studies for Willmore flow of sphere under spontaneous mean curvature for the solvers in List 7.2.
@@
Results are presented in Figure 9. The convergence in space and time is clearly visible and in accordance with the optimal expected. Moreover, we can see that the use of the mesh redistribution actually improves the convergence rates. We believe this behavior is due to the automatic mesh adaptivity of Equation 33 (already mentioned in the original article [36]), which tends to accumulate nodes in certain regions providing better geometry description.
As a second test we check the energy evolution for a torus of major radius $R=2$ and minor radius $r=1$ as done in [12, Section 5.1]. The minimizer of the Willmore energy in this case is the Clifford torus with radius $R=\sqrt{2},\;r=1$, which has energy $\mathcal{E}=4\pi^{2}\approx 39.4784$. We use both solvers and plot the energy history for two different examples described by the following parameter sets:
@@
The results are shown in Figure 9. We notice how both simulations using hfSolver 1 break before reaching the end time. This is due to instabilities that occur at regions of high curvature already observed in [12, Figure 2] for $\theta=1$. Applying the redistribution doesn’t impact the energy landscape and allows to proceed with the simulation until the stable state is reached. Moreover, refining the mesh shows how the energy converges toward the expected value.
@@ 7.2.1 Pinching tests for Helfrich system
We perform consolidated tests for the Willmore flow under the influence of spontaneous mean curvature as presented in [12, Section 5.3] and then repeated in [51], where a similar two-step procedure as the one adopted here is used. We consider two different tests for cigar-shaped surfaces:
@@
    hf-Test 1 with initial mesh as in 10(a) and the following sets of parameters
@@
    hf-Test 2 with initial mesh as in 10(d) and the following sets of parameters
@@
A common timestep of $\tau=10^{-3}$ has been used for all the experiments. The idea is to test the ability of the algorithm to simulate pinching phenomena, since hf-Test 1 develops one neck and hf-Test 2 develops two necks. The results are plotted in Figure 10 for both hfSolver 1 and 2. It can be seen how the mesh adaptivity mentioned in [36] leads to a better resolution of the neck and overall better results for hfSolver 2. By comparison of 10(c) with [12, Figure 12] the shape is qualitatively in accordance with what is expected. Even more, while [12, Figure 12] requires a finer mesh to accurately resolve the three pearls, the mesh adaptivity of [36] alleviates that requirement for similar resolution. Figure 10 is also in accordance with [51, Figure 5.14], where a similar algorithm is employed for a cigar-shaped surface with different aspect ratio. As it happens in that article, it has to be noted that the refinement of the necks comes to the expense of a coarser mesh at the lobes. Looking at the energy evolution for hf-Test 2, the curve for hfSolver 1 and hfSolver 2 fall in the same place for all experiments, highlighting that the tangential mesh redistribution does not impact the energy landscape of the shape dynamics. For hf-Test 1 we have that the dynamics is stable for both examples when $h=0.2$, with a slight difference in energy when necking occurs. Refining the mesh leads to mesh breaking for hfSolver 1, while hfSolver 2 maintains its stability, asymptotically tending towards the same energy hfSolver 1 was tending to in the coarser case. Overall, the proposed technique appears stable in various regimes and accurate in resolving the energy evolution of the geometry.

### v2

@@ 6.2 Numerical results for Cahn-Hilliard system
We proceed in reporting experimental studies to assess the accuracy of the proposed framework. Consider a cylinder revolving around the $x_{3}$-axis with radius $R=1$ and height $l=2$. To satisfy the assumption $\nabla_{\Gamma}\cdot{\boldsymbol{v}}=0$, we transform the initial cylinder through the isometry ${\boldsymbol{\Phi}}(t,{\boldsymbol{p}})={\boldsymbol{\Phi}}^{\mathcal{A}}(t,{\boldsymbol{p}})=\mathbb{A}(t){\boldsymbol{p}}+B(t)$
$\mathbb{A}(t)=\begin{bmatrix}\lambda(t)&0&0\\ 0&\lambda(t)&0\\ 0&0&1/\lambda(t)\end{bmatrix},\quad B(t)=\begin{bmatrix}0.2t\\ 0.1t\\ 0\end{bmatrix}$, (59)
@@
    chSolver 1: Solver (57) with no postprocessing.
@@
    chSolver 2: Solver (57) with imposed bound preservation in the interval (1,1)(-1,1).
@@
    chSolver 3: Solver (57) with imposed mass preservation using the lumped-mass inner product $(\cdot\,,\,\cdot)^{h}$.
@@
    chSolver 4: Solver (57) with imposed bound preservation in the interval (1,1)(-1,1) and mass preservation.
Zero flux Neumann boundary conditions are imposed for both $u$ and $w$. A random uniform initial distribution in the interval (1,1)(-1,1) is chosen for $u_{0}$. The parameters are set as
@@ 7 Membrane force system
In the realm of mathematical modeling, principal curvatures have long been used for the theory of elastic plates and shells, with the work of Kirchhoff Kirchhoff [1850] representing what may be the most famous case. Later in the ’70s, the work of Canham Canham [1970] and Helfrich Helfrich [1973] on the link between biological membranes shape and curvature functionals has de facto initiated an entirely new field of research focused on the use of geometry of hypersurfaces for biophysical modeling.
@@
$\mathcal{E}_{B}(t)=\int_{\Gamma}\frac{\gamma_{W}}{2}(\kappa-\kappa_{0})^{2}+\int_{\Gamma}\gamma_{G}\kappa_{G}+\gamma_{\partial}\int_{\partial\Gamma}1,$ (60)
where $\gamma_{W},\;\gamma_{G}\in\mathbb{R}$ are bending rigidities, $\kappa:\mathcal{G}_{T}\to\mathbb{R}$ is the previously introduced mean curvature, $\kappa_{0}\in\mathbb{R}$ is the spontaneous curvature, and $\kappa_{G}:\mathcal{G}_{T}\to\mathbb{R}$ is the Gaussian curvature $\kappa_{G}=\det\mathbb{W}$. The parameter $\gamma_{\partial}$ takes into account the possible line energy of $\partial\Gamma$. For the sake of simplicity we will consider the case $\gamma_{G}=constant$, $\gamma_{W}=constant$, and assume that no topological change occurs along the evolution. Taken into account these simplifications, the gradient flow of a membrane $\Gamma$ is a family of evolving surfaces with normal velocity $\boldsymbol{v}$$^{\perp}$ given by Nitsche [1993], Barrett et al. [2017a]
${\boldsymbol{v}}^{\perp}=\gamma_{W}\Bigl(-\Delta_{\Gamma}\kappa+\frac{1}{2}\kappa(\kappa-\kappa_{0})^{2}-(\kappa-\kappa_{0})|\mathbb{W}|^{2}\Bigr){\boldsymbol{n}}_{\Gamma}.$ (61)
@@
Remark 5.
Other geometrical flows like the mean curvature flow and the surface diffusion flow are also of interest in this field. Being somewhat simpler, extensive research is already available for them, while less is known about the Willmore flow, which we focus on in this article. We refer to Dziuk and Elliott [2007], Deckelnick et al. [2005], Barrett et al. [2020] for further details.
Remark 6.
@@ 7.1 Numerical methods for Helfrich system
We choose here the family of BGN-type algorithms as pioneered in Barrett et al. [2007] and later adapted for Willmore flow in Barrett et al. [2016, 2017a]. The schemes pivot around the two following ingredients:
  • Introduce the mean curvature vector ${\boldsymbol{\kappa}}(t)=\kappa(t){\boldsymbol{n}}_{\Gamma}(t)$ as independent variable and use the defining equation for the mean curvature vector ${\boldsymbol{\kappa}}=\Delta_{\Gamma}\mathbf{Id}_{\Gamma}$ in order to reduce the fourth order problem to a coupled second order one Barrett et al. [2007];
  • Use the weak formulation to rewrite and simplify the term $\kappa|\mathbb{W}|^{2}$ (more precisely $\kappa|\mathbb{H}|^{2}$ in the immersed setting) taking advantage of the newly introduced variable $\boldsymbol{\kappa}$}. This reformulation not only simplifies and linearizes the system but also leads to stable schemes as proven in Dziuk [2008].
Our scheme choice builds on the one derived in Barrett et al. [2017a] for boundary value problems. Among the different types of boundary conditions considered in Barrett et al. [2017a] we will restrict to the case of clamped boundary conditions, for which
@@
$({\boldsymbol{v}}^{\perp}\,,\,{\boldsymbol{\phi}})=~(\nabla_{\Gamma}{\boldsymbol{y}}\,,\,\nabla_{\Gamma}{\boldsymbol{\phi}})+(\nabla_{\Gamma}\cdot{\boldsymbol{y}}\,,\,\nabla_{\Gamma}\cdot{\boldsymbol{\phi}})-((\nabla_{\Gamma}{\boldsymbol{y}})^{T}\,,\,\mathcal{D}({\boldsymbol{\phi}})\;\mathbb{P}_{\Gamma})$ (63a)
@@
$({\boldsymbol{y}}\,,\,{\boldsymbol{\psi}})=~(\gamma_{W}({\boldsymbol{\kappa}}-\kappa_{0}{\boldsymbol{n}}_{\Gamma})\,,\,{\boldsymbol{\psi}}),$ (63b)
$({\boldsymbol{\kappa}}\,,\,{\boldsymbol{\xi}})=~-(\mathbb{P}_{\Gamma}\,,\,\nabla_{\Gamma}{\boldsymbol{\xi}})+({\boldsymbol{\mu}}\,,\,{\boldsymbol{\xi}})_{\partial\Gamma},$ (63c)
@@
${\boldsymbol{v}}_{h}^{\perp}(t_{n})=\frac{1}{\tau}\sum_{i}({\boldsymbol{y}}_{i}(t_{n+1})-{\boldsymbol{y}}_{i}(t_{n}))\phi_{i}(t_{n})=\frac{1}{\tau}{\boldsymbol{d}}_{h}^{n+1},$ (64)
where we recall that ${\boldsymbol{y}}_{i}$ are the mesh points and $\{\phi_{i}\}$ the associated basis such that $\phi_{i}({\boldsymbol{y}}_{j})=\delta_{ij}$ at all times. Moreover, (63c) and (63b) can be collapsed to
@@
$+(\kappa_{0}{\boldsymbol{n}}_{\Gamma^{n}}\,,\,{\boldsymbol{\psi}})_{\Gamma^{n}}+({\boldsymbol{\mu}}\,,\,{\boldsymbol{\psi}})_{\partial\Gamma^{n}}.$ (65)
@@
$\mathopen{}\left(\frac{{\boldsymbol{d}}_{h}^{n+1}}{\tau}\,,\,{\boldsymbol{\phi}}_{h}\mathclose{}\right)_{\Gamma_{h}^{n}}^{h}-(\nabla_{\Gamma_{h}^{n}}{\boldsymbol{y}}^{n+1}_{h}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\phi}}_{h})_{\Gamma_{h}^{n}}=~(\nabla_{\Gamma_{h}^{n}}\cdot{\boldsymbol{y}}^{n}_{h}\,,\,\nabla_{\Gamma_{h}^{n}}\cdot{\boldsymbol{\phi}}_{h})_{\Gamma_{h}^{n}}$ (66a)
@@
$\frac{1}{\gamma_{W}}({\boldsymbol{y}}^{n+1}_{h}\,,\,{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}^{h}+(\nabla_{\Gamma_{h}^{n}}{\boldsymbol{d}}_{h}^{n+1}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}$ (66b)
@@
Remark 8.
Note that (66) corresponds to the scheme presented in Barrett et al. [2016] for the choice of the parameter $\theta=1$, i.e. without tangential motion control.
@@ 7.2 Numerical results for Helfrich system
To test the accuracy of our method for gradient flows, we start by performing classical studies on Willmore flow such as the ones in Dziuk [2008], Barrett et al. [2016, 2017a]. For every test we compare
@@
    hfSolver 2: algorithm (66) together with the mesh redistribution described in (33). We highlight that it is of crucial importance that, when the algorithm in Section 4 is employed, geometrical quantities characteristic of the mesh are updated. In our case, we refer to the mean curvature ${\boldsymbol{\kappa}}^{n}_{h}$ and the auxiliary variable ${\boldsymbol{y}}^{n}_{h}$ of the algorithm in (66).
We start by testing the convergence using the example presented in [Barrett et al., 2016, Section 5.3]. It is an evolving sphere with initial radius $R_{0}$ and spontaneous curvature $\kappa_{0}$. The radius evolution $R(t)$ is the solution to the ordinary differential equation
$\partial_{t}R=-\frac{\kappa_{0}}{R}\Bigl(\frac{2}{R}+\kappa_{0}\Bigr),\quad R(0)=R_{0}>0.$ (67)
@@
$\lVert\Gamma-\Gamma_{h}\rVert_{L^{\infty}(L^{\infty})}=\max_{n}\max_{i}|\lVert{\boldsymbol{y}}_{i}(t_{n})\rVert-R(t_{n})|,$ (68)
@@
Figure 8: Convergence studies for Willmore flow of sphere under spontaneous mean curvature for the solvers in List 7.2. As expected, first and second order convergence are achieved in time and space, respectively.
@@
Results are presented in Figure 9. The convergence in space and time is clearly visible and in accordance with the optimal expected. Moreover, we can see that the use of the mesh redistribution actually improves the convergence rates. We believe this behavior is due to the automatic mesh adaptivity of (33) (already mentioned in the original article Duan and Li [2024]), which tends to accumulate nodes in certain regions providing better geometry description.
As a second test we check the energy evolution for a torus of major radius $R=2$ and minor radius $r=1$ as done in [Barrett et al., 2016, Section 5.1]. The minimizer of the Willmore energy in this case is the Clifford torus with radius $R=\sqrt{2},\;r=1$, which has energy $\mathcal{E}=4\pi^{2}\approx 39.4784$. We use both solvers and plot the energy history for two different examples described by the following parameter sets:
@@
The results are shown in Figure 9. We notice how both simulations using hfSolver 1 break before reaching the end time. This is due to instabilities that occur at regions of high curvature already observed in [Barrett et al., 2016, Figure 2] for $\theta=1$. Applying the redistribution doesn’t impact the energy landscape and allows to proceed with the simulation until the stable state is reached. Moreover, refining the mesh shows how the energy converges towards the expected value.
@@ 7.2.1 Pinching tests for Helfrich system
We perform consolidated tests for the Willmore flow under the influence of spontaneous mean curvature as presented in [Barrett et al., 2016, Section 5.3] and then repeated in Garcke et al. [2025], where a similar two-step procedure as the one adopted here is used. We consider two different tests for cigar-shaped surfaces:
@@
    hf-Test 1 with initial mesh as in Figure 10(a) and the following sets of parameters
@@
    hf-Test 2 with initial mesh as in Figure 10(d) and the following sets of parameters
@@
A common timestep of $\tau=10^{-3}$ has been used for all the experiments. The idea is to test the ability of the algorithm to simulate pinching phenomena, since hf-Test 1 develops one neck and hf-Test 2 develops two necks. The results are plotted in Figure 10 for both hfSolver 1 and 2. It can be seen how the mesh adaptivity mentioned in Duan and Li [2024] leads to a better resolution of the neck and overall better results for hfSolver 2. By comparison of Figure 10(c) with [Barrett et al., 2016, Figure 12] the shape is qualitatively in accordance with what is expected. Even more, while [Barrett et al., 2016, Figure 12] requires a finer mesh to accurately resolve the three pearls, the mesh adaptivity of Duan and Li [2024] alleviates that requirement for similar resolution. Figure 10 is also in accordance with [Garcke et al., 2025, Figure 5.14], where a similar algorithm is employed for a cigar-shaped surface with different aspect ratio. As it happens in that article, it has to be noted that the refinement of the necks comes to the expense of a coarser mesh at the lobes. Looking at the energy evolution for hf-Test 2, the curve for hfSolver 1 and hfSolver 2 fall in the same place for all experiments, highlighting that the tangential mesh redistribution does not impact the energy landscape of the shape dynamics. For hf-Test 1 we have that the dynamics is stable for both examples when $h=0.2$, with a slight difference in energy when necking occurs. Refining the mesh leads to mesh breaking for hfSolver 1, while hfSolver 2 maintains its stability, asymptotically tending towards the same energy hfSolver 1 was tending to in the coarser case. Overall, the proposed technique appears stable in various regimes and accurate in resolving the energy evolution of the geometry.